OpenStudy (anonymous):

-10+log base 3 (n+3)=-10

OpenStudy (gorv):

add 10 both side

OpenStudy (gorv):

\[\log_{3} (n+3)=0\]

OpenStudy (gorv):

n+3=3^0=1 n+3=1 subtract 3 from both side

OpenStudy (gorv):

n=1-3